# Using Large Language Models to Tackle Fundamental Challenges in Graph Learning: A Comprehensive Survey

### v1

@@ 3.3.2 Multimodal Attributed Graph Learning
(1) Traditional Methods for Multimodal Attribute Modeling Traditional methods aim to learn alignment or fusion patterns across different modalities by training models to capture the underlying relationships between heterogeneous multimodal attributes. MMGCN [211] learns a multimodal graph convolution network on a user-item bipartite graph to learn modal-specific representations of users and micro-videos to better capture user preferences in different modalities. scMoGNN [213] utilizes different GNNs to learn representations of each cell-feature bipartite graph in different modalities, and finally concatenates these representations to fuse information from different modalities. [214] and [215] introduce an attention mechanism in graph neural networks to dynamically capture the importance of information from different modalities. To reduce the need for large amounts of labeled data and improve robustness, contrastive learning methods are proposed for modeling multimodal graph data. Joyful [216] designs a global contextual fusion module and a specific modalities fusion module to capture information in different scales, and then concatenates the representations from both modules to generate a unified representation vector for each node. These modules are trained by comparing positive and negative pairs in the corrupted graphs through edge perturbation and random masking. FormNetV2 [217] proposes a centralized multimodal graph contrastive learning strategy to learn fused representations from different modalities in one loss. HGraph-CL [218] introduces a hierarchical graph contrastive learning framework that builds intra-modal and inter-modal graphs, leveraging graph augmentations and contrastive learning to capture complex sentiment relations within and across different modalities.
While these traditional methods can align or fuse multimodal attributes in graph data, they still rely on abundant data to train the model and cannot easily generalize to an unseen domain. Compared to LLMs, these traditional methods are less adaptable to diverse data distributions and struggle to leverage pre-trained knowledge for more efficient generalization.
@@
Recent research on MLLMs focuses on developing advanced methods to align LLMs with models from other modalities [219, 220, 221, 222, 223]. Although these methods are not specifically designed for graph data, they can be used as powerful preprocessing tools for aligning multimodal attributes and generating unified representations. For example, by unifying textual and visual side information through a pretrained model Clip-ViT [224], LLMRec [167] effectively mitigates the multimodal heterogeneity and enhances the node features in the recommendation network, leading to significant performance improvements. MAGB [21] conducts experiments on a set of large multimodal attributed graph datasets, which demonstrates that MLLMs can effectively alleviate the biases from cross-domain multimodal heterogeneity. Recent approaches align encoders for different modalities by leveraging graph structures, making them more naturally suited for graph learning tasks. MMGL [173] utilizes LLMs and image encoders with adapter layers to embed text and image attributes, respectively. These embeddings are then combined with graph positional encodings to capture graph structure information and finally fed into LLMs to generate the corresponding outputs. GraphAdapter [174] introduces GNN-based adaptors for encoders of different modalities, which can better align these encoders based on the graph structure information. TouchUp-G [175] improves node features obtained from pretrained models of different modalities by adapting them to the graph structure, using a new metric called feature homophily to quantify the correlation between the graph and node features, which enhances GNN performance across different tasks and data modalities. Following this direction, UniGraph2 [176] leverages modality-specific encoders alongside a GNN and an MoE module to effectively unify multimodal features while preserving the underlying graph structure.
@@ 3.3.3 Structural Heterogeneous Graph Learning
Beyond merely understanding topological structures, textual descriptions can also incorporate the attributes of nodes and edges, providing a richer context that enables LLMs to capture the intricate features of the graph structure. GraphText [183] constructs a graph-syntax tree to preserve the hierarchical structure and traverses the tree structure to generate a graph text sequence, where the node attributes are incorporated in the description of leaf nodes. WalkLM [184] generates paths by attributed random walks and textualizes the paths into natural languages for processing in LLMs. To effectively capture cross-group connections while minimizing noisy nodes, Path-LLM [185] employs shortest paths to generate structural sequences. InstructGLM [186] describes graph structures using neighbors in different scales and utilizes instruction tuning to finetune an LLM to better perform graph tasks. The textual node attributes are concatenated with the node index to better incorporate semantic information. MuseGraph [187] incorporates neighbors and paths together to textualize graph structures, which can capture both local connectivity and complex path-based relationships between nodes. Graph-LLM [188] transform text-attributed graphs into natural languages and demonstrates that LLMs can understand the graph structures from language descriptions and demonstrate remarkable zero-shot performance on graph tasks.
By converting diverse graph structures into natural language descriptions, these structure-to-text transformation methods enable LLMs to interpret and reason across varying connectivity patterns in a unified manner. These approaches allow LLMs to leverage their pre-trained linguistic knowledge to identify valuable relational patterns and facilitate more flexible reasoning across different graph topologies with excessive heterogeneity.
@@
GNP [189] employs a GNN to encode the knowledge graph structure into graph tokens and subsequently uses projectors to map these tokens into the text token space of an LLM. Since the projected graph tokens share the same embedding space as the original text tokens of the LLMs, they can be seamlessly concatenated with language instructions and fed into the LLM. This integration allows the LLM to leverage both graph structures and textual context to tackle various graph-related tasks. GraphToken [190] generates graph tokens through a GNN for general graphs and aligns the graph tokens with the token space of a frozen LLM. GraphGPT [191] introduces a dual-stage instruction tuning framework that directly aligns graph tokens with LLMs using both self-supervised and task-specific instruction tuning. LLaGA [192] reorganizes graph nodes into structure-aware sequences based on two different templates, then maps these sequences into the token embedding space of LLMs through a versatile projector, enabling the model to handle various graph tasks while maintaining the general-purpose nature of LLMs. TEA-GLM [193] utilizes feature-wise contrastive learning with token embeddings, where principal component analysis is applied to the token embeddings of an LLM to obtain principal components that serve as coordinate axes. These axes are used to map node representations generated from a GNN into the token space of the LLM. To generate graph tokens for heterogeneous graphs, HiGPT [194] introduces an in-context heterogeneous graph tokenizer that encodes diverse node and edge types by using a language-based parameterized heterogeneity projector, dynamically generating graph tokens that represent the semantic relationships in heterogeneous graphs.
@@ 3.3.4 Summary of Heterogeneity
LLMs, pre-trained on diverse datasets, are inherently capable of integrating information from various sources into a unified semantic space. The reviewed studies show that LLM-integrated models can transfer knowledge across different graph domains without the need for domain-specific retraining. By leveraging the generalizability of LLMs, these models are able to mitigate different forms of cross-domain heterogeneity, including textual attribute heterogeneity, multimodal attribute heterogeneity, and structural heterogeneity. As a result, they demonstrate superior performance in addressing cross-domain heterogeneity, particularly in scenarios where traditional methods are ineffective or inapplicable due to the lack of sufficient training data. Experiments in these papers show improved performance on cross-domain node classification, link prediction, and graph classification tasks when LLMs are used to either encode heterogeneous content or assist in aligning representations.
@@ 3.4 Dynamic Instability in Graphs
Figure 12: Illustration of dynamic instability in graphs. The figure presents three categories of dynamic instability over two timestamps ($t_{1}$rightarrow $t_{2}$): (1) Node Feature Dynamics, where node attributes change over time; (2) Structural Dynamics, showing newly added node (solid dark purple circle) and newly removed node (gray dashed circle); (3) New Node/Edge Types, depicting the emergence of entirely new node and relationship types, represented by a solid dark purple triangle and dashed edges.
@@
Table 4: Literature Summary of LLM-based Methods for Handling Dynamic Instability in Graphs.
Domain Type Reference (Method—Citation) Querying and Reasoning Forecasting & Reasoning ICL[235], zrLLM[236], CoH[237], TG-LLM[238], LLM4DyG[239] QA & Interpretability TimeR4[240], GenTKGQA[241], Unveiling LLMs[242] Generating and Updating Generating Structures FinDKG[243], GenTKG[244], Up To Date[245], PPT[246], LLM-DA[247] Generating Insights & Representations TimeLlama[248], RealTCD[249], DynLLM[250] Evaluation and Application Model Evaluation Dynamic-TempLAMA[251], DARG[252] Downstream Applications AnomalyLLM[83], MoMa-LLM[253], TRR[254]
@@ 3.4.3 Evaluation and Application in Dynamic Graphs
(2) Model Evaluation LLMs provide new perspectives for evaluating model performance and robustness in dynamic settings. Dynamic-TempLAMA [251] presents a dynamic benchmarking framework specifically designed to assess how well pre-trained language models (MLMs) handle temporal concept drift—the evolution of factual knowledge over time. The framework dynamically creates time-sensitive test sets from Wikidata and evaluates MLMs through multiple views (probing, generation, scoring) to determine if their internal knowledge is outdated. DARG [252] addresses the limitations of static benchmarks by proposing a method to dynamically generate new evaluation data. It introduces changes to the reasoning graphs of existing benchmark samples to create novel test data with controlled complexity and diversity, using a code-augmented LLM to ensure label correctness. This enables adaptive evaluation of LLMs’ reasoning capabilities as they evolve. Both works highlight the importance of moving beyond static, snapshot-based evaluations toward more dynamic, adaptive approaches, especially as both models and world knowledge constantly change.

### v2

@@ 3.3.2 Multimodal Attributed Graph Learning
(1) Traditional Methods for Multimodal Attribute Modeling Traditional methods aim to learn alignment or fusion patterns across different modalities by training models to capture the underlying relationships between heterogeneous multimodal attributes. MMGCN [289] learns a multimodal graph convolution network on a user-item bipartite graph to learn modal-specific representations of users and micro-videos to better capture user preferences in different modalities. scMoGNN [291] utilizes different GNNs to learn representations of each cell-feature bipartite graph in different modalities, and finally concatenates these representations to fuse information from different modalities. [292] and [293] introduce an attention mechanism in graph neural networks to dynamically capture the importance of information from different modalities. To reduce the need for large amounts of labeled data and improve robustness, contrastive learning methods are proposed for modeling multimodal graph data. Joyful [294] designs a global contextual fusion module and a specific modalities fusion module to capture information at different scales, and then concatenates the representations from both modules to generate a unified representation vector for each node. These modules are trained by comparing positive and negative pairs in the corrupted graphs through edge perturbation and random masking. FormNetV2 [295] proposes a centralized multimodal graph contrastive learning strategy to learn fused representations from different modalities in one loss. HGraph-CL [296] introduces a hierarchical graph contrastive learning framework that builds intra-modal and inter-modal graphs, leveraging graph augmentations and contrastive learning to capture complex sentiment relations within and across different modalities.
While these traditional methods can align or fuse multimodal attributes in graph data, they still rely on abundant data to train the model and cannot easily generalize to an unseen domain. Compared to LLMs, these traditional methods are less adaptable to diverse data distributions and struggle to leverage pretrained knowledge for more efficient generalization.
@@
Recent research on MLLMs focuses on developing advanced methods to align LLMs with models from other modalities [297, 298, 299, 300, 301]. Although these methods are not specifically designed for graph data, they can be used as powerful preprocessing tools for aligning multimodal attributes and generating unified representations. For example, by unifying textual and visual side information through a pretrained model Clip-ViT [302], LLMRec [217] effectively mitigates the multimodal heterogeneity and enhances the node features in the recommendation network, leading to significant performance improvements. MAGB [24] conducts experiments on a set of large multimodal attributed graph datasets, which demonstrate that MLLMs can effectively alleviate the biases from cross-domain multimodal heterogeneity. Recent approaches align encoders for different modalities by leveraging graph structures, making them more naturally suited for graph learning tasks. MMGL [230] utilizes LLMs and image encoders with adapter layers to embed text and image attributes, respectively. These embeddings are then combined with graph positional encodings to capture graph structure information and finally fed into LLMs to generate the corresponding outputs. GraphAdapter [232] introduces GNN-based adaptors for encoders of different modalities, which can better align these encoders based on the graph structure information. TouchUp-G [244] improves node features obtained from pretrained models of different modalities by adapting them to the graph structure, using a new metric called feature homophily to quantify the correlation between the graph and node features, which enhances GNN performance across different tasks and data modalities. Following this direction, UniGraph2 [245] leverages modality-specific encoders alongside a GNN and an MoE module to effectively unify multimodal features while preserving the underlying graph structure.
@@ 3.3.3 Structural Heterogeneous Graph Learning
Beyond merely understanding topological structures, textual descriptions can also incorporate the attributes of nodes and edges, providing a richer context that enables LLMs to capture the intricate features of the graph structure. GraphText [252] constructs a graph-syntax tree to preserve the hierarchical structure and traverses the tree structure to generate a graph text sequence, where the node attributes are incorporated in the description of leaf nodes. WalkLM [254] generates paths by attributed random walks and textualizes the paths into natural languages for processing in LLMs. To effectively capture cross-group connections while minimizing noisy nodes, Path-LLM [255] employs shortest paths to generate structural sequences. InstructGLM [256] describes graph structures using neighbors in different scales and utilizes instruction tuning to finetune an LLM to better perform graph tasks. The textual node attributes are concatenated with the node index to better incorporate semantic information. MuseGraph [257] incorporates neighbors and paths together to textualize graph structures, which can capture both local connectivity and complex path-based relationships between nodes. Graph-LLM [260] transforms text-attributed graphs into natural languages and demonstrates that LLMs can understand the graph structures from language descriptions and demonstrate remarkable zero-shot performance on graph tasks.
By converting diverse graph structures into natural language descriptions, these structure-to-text transformation methods enable LLMs to interpret and reason across varying connectivity patterns in a unified manner. These approaches allow LLMs to leverage their pretrained linguistic knowledge to identify valuable relational patterns and facilitate more flexible reasoning across different graph topologies with excessive heterogeneity.
@@
Advanced methods bridge the gap between graph structures and LLMs by converting graph data into graph tokens that LLMs can understand. GNP [261] and GraphToken [268] both use a GNN to encode graph structure and then employ a projector to map these embeddings into the token space of the LLM, allowing them to be processed alongside regular text. LLaGA [270] takes a different approach by reorganizing graph nodes into structure-aware sequences before projecting them, which helps the model maintain its general-purpose nature. TEA-GLM [271] refines this alignment with feature-wise contrastive learning, using principal components from the token space to precisely map GNN representations. GraphGPT [269] introduces a dual-stage instruction tuning framework that directly aligns graph tokens with LLMs using both self-supervised and task-specific instruction tuning. To generate graph tokens for heterogeneous graphs, HiGPT [272] introduces an in-context heterogeneous graph tokenizer that encodes diverse node and edge types by using a language-based parameterized heterogeneity projector, dynamically generating graph tokens that represent the heterogeneous semantic relationships.
@@
3.3.4 Evaluation for Handling Cross-Domain Heterogeneity in Graphs
We summarize the existing evaluation settings for LLM-based methods for handling cross-domain heterogeneity in graphs, covering benchmark datasets, evaluation metrics, and downstream tasks (see Table 3).
Evaluations typically use well-established traditional graph benchmark datasets. These datasets span a wide range of domains, including citation networks (e.g., Cora [81], Citeseer [85], and PubMed [82]), biomedical graphs (e.g., MIMIC-III [221], MOLHIV [224], and MOLPCBA [224]), commercial and recommendation networks (e.g., MovieLens [218], Netflix [219], and various Amazon product datasets [244]), as well as knowledge graphs (e.g., FB15K237 [94] and WN18RR [124]). While most of these datasets are standard in graph learning, cross-domain evaluations differ from conventional setups. Instead of training and testing within the same domain, LLM-based methods are often trained on source-domain datasets (or without further training) and tested on unseen target-domain datasets [252, 250, 227]. In some cases, few-shot settings are also utilized to assess generalization ability when only a limited number of labeled examples in the target domain are available [226, 213].
The choice of evaluation metrics typically depends on the task type. Classification tasks adopt Accuracy, Macro-F1, Micro-F1, or AUC, ranking-oriented tasks like recommendation or link prediction use MRR, NDCG, or Hits@N, and generation tasks such as graph-to-text generation rely on BLEU-4, ROUGE-L, CIDEr, METEOR, or CHRF++. Metrics are usually evaluated under zero-shot or few-shot settings to reflect cross-domain generalization. However, traditional metrics may miss cases where outputs are semantically correct but inconsistent in domain-specific format or style. LLMs, in contrast, can capture nuanced correctness across domains based on their strong language comprehension ability, positioning LLM-as-a-Judge [310] a promising direction for future evaluation. The downstream tasks range from node-level and edge-level tasks to graph-level tasks, including node classification, link prediction, and graph classification, as well as more complex settings like recommendation, question answering, section summarization, and graph-to-text generation. This broad coverage sufficiently evaluates the cross-domain generalization ability of LLM-based methods across diverse task types and application scenarios.
@@ 3.3.5 Summary of Heterogeneity
LLMs, pretrained on diverse datasets, are inherently capable of integrating information from various sources into a unified semantic space. The reviewed studies show that LLM-integrated models can transfer knowledge across different graph domains without the need for domain-specific retraining. By leveraging the generalizability of LLMs, these models are able to mitigate different forms of cross-domain heterogeneity, including textual attribute heterogeneity, multimodal attribute heterogeneity, and structural heterogeneity. As a result, they demonstrate superior performance in addressing cross-domain heterogeneity, particularly in scenarios where traditional methods are ineffective or inapplicable due to the lack of sufficient training data. Experiments in these papers show improved performance on cross-domain node classification, link prediction, and graph classification tasks when LLMs are used to either encode heterogeneous content or assist in aligning representations.
@@ 3.4 Dynamic Instability in Graphs
Figure 12: Illustration of dynamic instability in graphs. The figure presents three categories of dynamic instability over two timestamps ($t_{1}$ $\rightarrow$ $t_{2}$): (1) Node Feature Dynamics, where node attributes change over time; (2) Structural Dynamics, showing newly added node (solid dark purple circle) and newly removed node (gray dashed circle); (3) New Node/Edge Types, depicting the emergence of entirely new node and relationship types, represented by a solid dark purple triangle and dashed edges.
@@
Table 4: LLM-based methods for handling dynamic instability in graphs, grouped by tasks, with representative methods, datasets, metrics, and downstream tasks.
Domain Category Method Typical Datasets Common Metrics Downstream Tasks Querying and Reasoning Forecasting & Reasoning ICL [314] WIKI [315], YAGO [316], ICEWS14 [126], ICEWS18 [317] MRR, Hits@N Link Prediction zrLLM [318] ICEWS [319], ACLED [320] MRR, Hits@N Link Prediction CoH [321] ICEWS14 [126], ICEWS18 [317], ICEWS05-15 [126] MRR, Hits@N Link Prediction TG-LLM [322] TGQA [323], TimeQA [324], TempReason [325] F1, Accuracy, Exact Match Temporal Reasoning LLM4DyG [326] Enron [327], DBLP [328], Flights [329] Accuracy, F1, Recall Spatial-Temporal Reasoning, Graph Reasoning and Querying, Link Prediction QA & Interpretability TimeR4 [330] MULTITQ [331], TimeQuestions [332] Hits@N Temporal Knowledge Graph Question Answering GenTKGQA [333] CronQuestion [334], TimeQuestions [332] Hits@N Temporal Knowledge Graph Question Answering Unveiling LLMs [335] FEVER [336], CLIMATE-FEVER [337] Precision, Recall, F1, ROC AUC, Accuracy Claim Verification Generating and Updating Generating Structures FinDKG [338] WIKI [315], YAGO [316], ICEWS14 [126] MRR, Hits@N Link Prediction GenTKG [339] ICEWS14 [126], ICEWS18 [317], GDELT [340], YAGO [316] Hits@N Link Prediction Up To Date [341] Wikidata [129] Accuracy, Response Rate Fact Validation, Question Answering PPT [342] ICEWS14 [126], ICEWS18 [317], ICEWS05-15 [126] MRR, Hits@N Link Prediction LLM-DA [343] ICEWS14 [126], ICEWS05-15 [126] MRR, Hits@N Link Prediction Generating Insights & Representations TimeLlama [344] ICEWS14 [126], ICEWS18 [317], ICEWS05-15 [126] Precision, Recall, F1, BLEU, ROUGE Event Forecasting, Explanation Generation RealTCD [345] Simulation Datasets [345] Structural Hamming Distance, Structural Interventional Distance Temporal Causal Discovery, Anomaly Detection DynLLM [346] Tmall [347], Alibaba [348] Recall@K, NDCG@K Dynamic Graph Recommendation, Top-K Recommendation Evaluation and Application Model Evaluation Dynamic-TempLAMA [43] DYNAMICTEMPLAMA [349] Accuracy, MRR, ROUGE, F1 Temporal Robustness Evaluation, Factual Knowledge Probing DARG [80] GSM8K [350], BBQ [351], BBH Navigate [352], BBH Dyck Language [352] Accuracy, Complexity-Induced Accuracy Retention Rate, Exact Match, Accuracy Mathematical Reasoning, Social Reasoning, Spatial Reasoning, Symbolic Reasoning Downstream Applications AnomalyLLM [102] UCI Messages [103], Blogcatalog [104] AUC Anomaly Detection MoMa-LLM [353] iGibson scenes [354] AUC, Recall Semantic Interactive Object Search TRR [355] Reuters Financial News [356] AUROC Event Detection
@@ 3.4.3 Evaluation and Application in Dynamic Graphs
(2) Model Evaluation LLMs provide new perspectives for evaluating model performance and robustness in dynamic settings. Dynamic-TempLAMA [43] presents a dynamic benchmarking framework specifically designed to assess how well pretrained language models (MLMs) handle temporal concept drift—the evolution of factual knowledge over time. The framework dynamically creates time-sensitive test sets from Wikidata and evaluates MLMs through multiple views (probing, generation, scoring) to determine if their internal knowledge is outdated. DARG [80] addresses the limitations of static benchmarks by proposing a method to dynamically generate new evaluation data. It introduces changes to the reasoning graphs of existing benchmark samples to create novel test data with controlled complexity and diversity, using a code-augmented LLM to ensure label correctness. This enables adaptive evaluation of LLMs’ reasoning capabilities as they evolve. Both works highlight the importance of moving beyond static, snapshot-based evaluations toward more dynamic, adaptive approaches, especially as both models and world knowledge constantly change.
@@
3.4.4 Evaluation for Handling Dynamic Instability in Graphs
We review the evaluation pipeline of LLM-based methods for addressing dynamic instability in graphs, covering benchmark datasets, evaluation metrics, and downstream tasks (see Table 4). Commonly used datasets include temporal knowledge graph benchmarks such as ICEWS14 [126], ICEWS18 [317], ICEWS05-15 [126], YAGO [316], and WIKI [315] for link prediction and temporal reasoning. Event-based corpora such as ACLED [320], GDELT [340], and TempReason [325] are used for temporal forecasting, while domain-specific datasets like Enron [327], DBLP [328], Tmall [347], and Alibaba [348] support communication analysis and recommendation. Temporal KGQA relies on MULTITQ [331], TimeQuestions [332], and CronQuestion [334], and claim verification uses FEVER [336] and CLIMATE-FEVER [337]. Anomaly and event detection adopt UCI Messages, Blogcatalog, iGibson scenes, and Reuters Financial News.
Evaluation metrics vary by task: MRR and Hits@N dominate ranking tasks; Accuracy, F1, Precision, Recall, and ROC-AUC are applied in classification; BLEU, ROUGE, and Exact Match assess generative tasks; Recall@K and NDCG@K measure recommendation; and SHD or SID evaluate temporal causal discovery. AUC or AUROC are common for anomaly detection.
The surveyed methods span diverse downstream tasks, including link prediction, temporal reasoning, temporal KGQA, claim verification, fact validation, dynamic graph recommendation, temporal causal discovery, event forecasting, and anomaly or event detection. This diversity highlights the capability of LLMs to integrate temporal reasoning and external knowledge for robust performance in evolving graph environments.